What is 2:3 ratio


1
Expert's answer
2020-09-10T17:12:23-0400

Ratio indicates how many times one number or thing contains another.

Let's explain what 2:3 ratio is:

For example we have eight(8) red balls and twelve(12) green balls in one box.

Find greatest common divisor of 8 and 12:

8 = 2 * 2 * 2 and 12 = 2 * 2 * 3

So their common divisors are 2 * 2:

8 = 2 * 2 * 2 and 12 = 2 * 2 * 3,

then denote 2 * 2 as X:

8 = X * 2 and 12 = X * 3

We can see that 8 and 12 correlate as 2 and 3.

So ratio between 8 and 12 is 2 : 3 and ratio of red balls and green balls in one box is also 2 : 3.


Another example.

Alex has a pool which size is 100 m х 150 m(Picture 1).



If we if we divide into equal segments(greatest common divisor of 100 and 150 is 50):


so we also get that ratio between 100 and 150 is 2 : 3. We can say that ratio between width and length of the pool is 2 : 3.